Given the sequence 20, 24, 28, 32, 36, . . . find the 60th term.
a. 216 b. 64 c. 256 d. 60 e. 80

Answer by stanbon(75887)

Given the sequence 20, 24, 28, 32, 36, . . . find the 60th term.
------
nth term pattern: a(n) = 20 + (n-1)4
60th term: a(60) = 20 + 59*4 = 256
=======================================
Cheers,
Stan H.
========================================
a. 216 b. 64 c. 256 d. 60 e. 80